Write
z1 and z2 in polar form. (Express θ in radians. Let 0 ≤ θ < 2π.)
z1 = sqrt(3)+ i, z2 = 1 + sqrt(3i)
Find
z1z2
z1/z2
1/z1

Answer by stanbon(75887)

Write
z1 and z2 in polar form. (Express θ in radians. Let 0 ≤ θ < 2π.)
z1 = sqrt(3)+ i,
Note:: z1 is in QI
-----
r1 = sqrt((sqrt(3))^2+1^2) = 2
theta = arctan(1/sqrt(3)) = pi/6
So r1 = 2cis(pi/6)
=========================================
z2 = 1 + sqrt(3)i
r2 = 2
theta = arctan(sqrt(3)) = pi/3
So r2 = 2cis(pi/3)
------

Find
z1z2 = r1*r2*cis(pi/6+2pi/6) = 4*cis(pi/2)
z1/z2 = r1/r2*cis(pi/6-2pi/6) = cis(-pi/6)
1/z1 = (1+ 0i)/((sqrt(3)+i) = cis(0)/[2*cis(pi/6)] = (1/2)cis(-pi/6)
------------
Cheers,
Stan H.
-----------